WIYN Observatory

WIYN Operational Status

Below is the current status of known issues that potentially impact operations or instrument readiness at WIYN. Questions about any of these items can be directed to Jayadev Rajagopal (jayadev.rajagopal at @gmail.com).

Telescope

No known issues.

ODI

ODI shutter closed when humidity ≥ 80% [02/28/2017]
The ODI dewar window is prone to condensation under high humidity conditions. Since condensation will leave a residue, and every condensation event will fog up some parts of the dewar window long-term. This diminishes the throughput significantly (up to a factor of 2), and in a patchy, non-calibratable way. It is therefore imperative to avoid any condensation event.

Until further notice, ODI observations shall pause (i.e., do not open the ODI shutter for an exposure) while the WIYN-reported humidity is ≥ 70%. Note: this only applies to the ODI shutter.

UPDATE! [02/28/2017] Based on positive results from increasing ODI's operational humidity limit in 2016B, we are increasing the operational limit for the remainder of 2017A from 70% RH to 80% RH. This increase is on a trial basis and will be continuously reassessed.

HYDRA

Fibers crossing [06/28/2016]
Hydra has seen an increase in fibers crossing. The issue is believed to be due to the age of the system. The zone of avoidance radius has been increased as of mid-June. Observers are cautioned to re-check their Hydra configurations with the updated simulator. Please contact your instrument support person for further information.

2 fibers removed from service [06/28/2016]
Two additional fibers have been removed from service due to potential damage - numbers 41 and 86. Please use the updated concentricities file when making your configurations. Contact your instrument support person for more information.

FOP 24 removed from concentricities [02/19/2016]
FOP 24 has been removed from the current concentricities files after tests have shown lower than expected throughput on that FOP. More testing is needed to understand the cause. For now we are removing the FOP from service in order to restrict handling and potential damage.

WHIRC

No known issues.

WTTM continues to work well and should be used as conditions permit.

Bench Spectrograph

No known issues.

IFUs (SPSK/HexPK/GradPK)

No known issues

Last modified: 28-Feb-2017 16:36:14 MST